OpenStudy (anonymous):

College Algebra: Anyone good with word problems? I'm doing 2x2 Linear Systems and I need to solve this problem using two variables and a system of two equations. I can use any method: Substitution or Addition. This is the problem: "Two-Income Family: Althea has a higher income than Vaughn and their total income is $82,000. If their salaries differ by $16,000, then what is the income of each?"

OpenStudy (freckles):

So let's call Althea's income A. And let's call Vaughn's income V. It says their total income is 82000. Can you translate that into a math equation?

OpenStudy (freckles):

And we are given Althea's income is higher than Vaughn's. And we know the incomes differ by 16000. So you should ale be able to come up with a math equation for this.

OpenStudy (dan815):

x+y=82000 x-y=16000

OpenStudy (freckles):

When I think of total, I think of sum or addition. When I think of differ by, I think of difference or subtraction. Or dan can just give you the equations...

OpenStudy (dan815):

:>

OpenStudy (anonymous):

lmao nice Dan. Alright, so since it's trying to figure out the income of each of them, why would it be x+y and x-y?

OpenStudy (dan815):

let x be money earned by althea let y be money earned by vaughn x+y=82000 x-y=16000

OpenStudy (anonymous):

Okay, But what do I add to that equation? How can I solve it if those are the only two numbers the problem gave?

OpenStudy (dan815):

if x-y=16000 then x=16000+y x+y=82000 16000+y+y=82000

OpenStudy (anonymous):

So, (16,000+y)+y=82,000. Wouldn't that be 16,001y+y=82,000. Then 16,002y=82,000.

OpenStudy (dan815):